Mr. Lindseyhad to order some new risers for the upcoming choral concert.The total for the small, lower tier is one dollar less than the upper, larger tier. Since it costs money from the cjoir budget, the numbers have to be entered as "negative numbers. The product of two consecutive negative integers as 1122. What are the numbers?
Answer by oberobic(2304)

Two consecutive negative integers are: x and x+1.
.
x*(x+1) = 1122
x^2 +x -1122 = 0
.
To factor this equation you need two value to multiply that are 1 apart and = 1122.
Given the small difference, the square root is a good first guess.
sqrt(1122)=33.4962
.
So test 33 and 34.
33 * 34 = 1122
.
(x-33)(x+34) = 0
.
So, x = 33 or -34.
We're looking for negative numbers, so x = -34.
x+1 = -33
.
Answer: The two numbers are -33 and -34.
.
Done.